OpenStudy (anonymous):

the area of a rectangle is 54 cm^2. The side lengths are 2x +1 and x+2. What is the measure of each side?

hartnn (hartnn):

Area of rectangle in terms of side is ?

OpenStudy (anonymous):

A=bh so 54+ (2x+1) (x+2) i got that much,,

hartnn (hartnn):

54=(2x+1) (x+2) now foil (2x+1) (x+2)

OpenStudy (anonymous):

2x^2 +4x +x +2 =54

OpenStudy (anonymous):

then 2x^2 +5x +2 +54

OpenStudy (anonymous):

2x^2 +5x -52 =0

OpenStudy (anonymous):

is that right??

OpenStudy (anonymous):

that is as far as i can get

hartnn (hartnn):

yes, that is correct. now tell me factors of 52

hartnn (hartnn):

not 52,of 52*2 = 104

OpenStudy (anonymous):

2x2 x 13

hartnn (hartnn):

ok, so of 104 it will be 2x2x2 x 13 = 8x13 now can u factorize ? write 5x as 13x-8x

OpenStudy (anonymous):

i don't get this part

hartnn (hartnn):

2x^2 +5x -52 =0 2x^2 +8x -3x -52 = 0 we basically need 2 numbers whose product is -52x2 =-104 and sum is 5. so we need to factor 104 as 8x13 so we get those 2 numbers as 8 and -3 (sum=5) so we write 5x = 8x-3x now u understand ?

OpenStudy (anonymous):

i have to think about that

hartnn (hartnn):

sorry there is a typo, 104 = 8x13 so we write 5x as 13x-8x (and not 3x)

hartnn (hartnn):

2 numbers are -8 and 13 product =-104 , sum =5

OpenStudy (anonymous):

so x =4

OpenStudy (anonymous):

this stuff is way beyond my understanding

hartnn (hartnn):

yes, x=4 is correct. the 2 values of x u get is 4,-13/2 but measure of side cannot be negative. so x is only 4 now put x=4 in 2x+1 and x+2 to get the sides

OpenStudy (anonymous):

9cm and 6cm,, that part is easy

OpenStudy (anonymous):

i guess i don't understand the factoring

hartnn (hartnn):

let me show you steps: 2x^2 +5x -52 =0 2x^2 -8x+13x -52 = 0 2x(x-4) +13 (x-4) = 0 (2x+13) (x-4 ) = 0 2x+13 = 0 OR x-4 =0 x= 13/2 OR x=4 when we have ax^2+bx+c=0 we find 2 numbers whose product is c*a and sum is b

OpenStudy (anonymous):

ok i'll study this,,,, tyty!

hartnn (hartnn):

welcome ^_^ do you want a simple example which you can try to factor ?

OpenStudy (anonymous):

yes please

hartnn (hartnn):

x^2 + 12x + 20 = 0

hartnn (hartnn):

2 numbers, product 20 , sum 12 ..... ?

OpenStudy (anonymous):

oh bother

hartnn (hartnn):

factor 20

OpenStudy (anonymous):

ohhh wait

OpenStudy (anonymous):

2x2x5

hartnn (hartnn):

so now can u find 2 numbers with sum = 12 using 2x2x5 = 20

OpenStudy (anonymous):

2 and 10????

hartnn (hartnn):

that is correct! good :) so write 12 x as 10x+2x can u continue now ? x^2 +10 x+ 2x +20 = 0

OpenStudy (anonymous):

i think so

hartnn (hartnn):

so tell me both factors?

OpenStudy (anonymous):

20 and 10

hartnn (hartnn):

i meant factors of x^2 + 12x + 20

hartnn (hartnn):

or could you work out values of x in x^2 + 12x + 20 =0

OpenStudy (anonymous):

x^2 (10x) (2x) +20 +0

OpenStudy (anonymous):

(x*X) (10x) (2x) +20 +0

hartnn (hartnn):

write it properly as x*x+ 10x +2x +20 = 0 now factor out x from first 2 terms and 2 from other 2 terms, what u get ?

OpenStudy (anonymous):

14x+20+0

hartnn (hartnn):

how u got 14 ??

hartnn (hartnn):

x*x is not 2x

OpenStudy (anonymous):

i added the x's

OpenStudy (anonymous):

12x +20=0

hartnn (hartnn):

x*x+ 10x +2x +20 = 0 x(x+10) + 2 (x+10) = 0 did u understand this step ?

OpenStudy (anonymous):

yes, i think so

hartnn (hartnn):

next step : (x+10)(x+2) =0 this ?

OpenStudy (anonymous):

yip,, that one i get

hartnn (hartnn):

next step : x+10 =0 OR x+2 = 0

OpenStudy (anonymous):

don't know

hartnn (hartnn):

when (x+a)(x+b) =0 then either of x+a or x+b has to be 0 , so that their product is 0 hence x+10 =0 OR x+2 = 0 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that makes sense

hartnn (hartnn):

now can u find 2 values of x from x+10 =0 OR x+2 = 0 ?

OpenStudy (anonymous):

yes